OpenStudy (anonymous):

7(3.5x10^4) in scienific notation

OpenStudy (anonymous):

get the product first of 7 and 3.5...

OpenStudy (triciaal):

standard form for scientific notation is a * 10^n where a is a number between 0 and 10 and n is an integer simply move the decimal to put any number in this format. To move to the left you are dividing by 10 the number of places is the exponent to keep the balance multiply by 10 to the opposite (that is the negative) of that exponent. To move to the right the reverse is true, example 0.25 = 2.5 * 10 ^ -1 example 358.2 = 3.582 * 10 ^2